Suppose a chemistry student can obtain two acid solutions from the stock room. The first solution is 20% acid and second solution is 45% acid. How many millimeters of each solution should the student mix together to obtain 100 millimeters of 30% acid solution?
Answer by stanbon(75887)

Suppose a chemistry student can obtain two acid solutions from the stock room. The first solution is 20% acid and second solution is 45% acid. How many millimeters of each solution should the student mix together to obtain 100 millimeters of 30% acid solution?
------
Equation:
acid + acid = acid
0.20x + 0.45(100-x) = 0.30*100
----
20x + 45*100 - 45x = 30*100
-25x = -15*100
x = 60 ml (amt of 20% solution needed)
100-x = 40 ml (amt of 45% solution needed)
===============================================
Cheers,
Stan H.